# Gravitational wave propagation in bigravity in the late universe

### v1

@@ Abstract
We carry out a detailed analytical investigation of the propagation of gravitational waves in bimetric gravity in a late-time de Sitter epoch. In this regime, the dynamical equations for the massless and massive graviton modes can be decoupled and solved exactly. We provide uniform approximations for the modes in terms of elementary functions, which are valid on all scales and for all viable mass windows. We identify different dynamical regimes for the system, depending on the propagation properties of the massive graviton, and whether the massless and massive components of the signal can be temporally resolved or not. In each regime, we compute the gravitational-wave luminosity distance as a function of redshift and study the propagation of wave packets. Further, by an explicit computation, we show that the massless and massive components of the signal retain their coherence also in the regime where they can be temporally resolved, even when couplings to incoherent matter degrees of freedom are included.
@@ 1 Introduction
A particularly relevant example is bimetric gravity [7, 8], which extends GR by considering two independent, dynamical and nonlinearly interacting metric fields. The interactions between the two metrics are constructed in a specific way that ensures the absence of the Boulware–Deser ghost [8, 9]. To maintain ghost-freedom, matter fields are coupled to only one of the two metrics [10, 11]. The theory can also be interpreted as describing the nonlinear dynamics of two interacting spin-2 fields (gravitons), one massless and the other massive, carrying 7 dynamical degrees of freedom in total. The free parameters associated with the interaction terms in the bimetric action can be rearranged to a set of 5 physical parameters: the effective energy density of dark energy $\Omega_{\Lambda}$, the mixing angle $\theta$ between the massless and massive gravitons, the mass $m_{\rm FP}$ of the massive graviton, and two additional parameters that determine the screening mechanism on solar system scales [12, 13]. Bimetric gravity admits two important limiting cases: GR is recovered in the $\theta\to 0$ limit, whereas massive gravity is obtained in the limit $\theta\to\frac{\pi}{2}$. The theory is compatible with all observational tests performed so far [14, 15, 13, 16], and has the potential to alleviate cosmological tensions [17, 18, 19, 20].
@@
In bimetric gravity, tensor perturbations are produced in states that are linear superpositions of massless and massive gravitons. Therefore, as they propagate, they undergo a mixing phenomenon that displays some analogies with neutrino and neutral-kaon oscillations. This was first noticed in Ref. [21] for proportional backgrounds [^1] That is, backgrounds such that the two metrics are conformally related with constant conformal factor, $f_{ab}=C^{2}\,g_{ab}$. For a cosmological background, this is equivalent to having a constant ratio of the scale factors and same conformal times., and later revisited in Refs. [22, 23] for cosmological backgrounds. As discussed in Ref. [22], in the regime where the gravitational wave signal component carried by the massive graviton cannot be temporally resolved from the massless component, the presence of the massive graviton induces a modulation of the gravitational wave signal, referred to as ‘gravitational wave oscillations’. Moreover, for gravitational wave signals with limited extent (wave packets) and sufficiently large propagation distances, the difference in group velocities between the massless and massive modes leads to a separation of the corresponding wave packets [23]. This has two main effects: a constant suppression factor in the amplitude of the GR-like signal carried by the massless component, and the existence of a secondary signal, interpreted as a distorted ‘echo’, carried by the massive mode. Previous analyses have placed constraints on the theory using gravitational wave observations, leaving some windows open corresponding to mixing angles $\theta\lesssim 10^{\circ}$. Specifically, the viable regions of parameter space, which are also compatible with all the other theoretical and observational constraints obtained so far [23, 12, 13, 16, 24], include (in units of the Hubble constant $H_{0}$): a low-mass range $1\lesssim m_{\rm FP}/H_{0}\lesssim 10^{4}$; an intermediate-mass range with $10^{6}\lesssim m_{\rm FP}/H_{0}\lesssim 10^{8}$ ; and a large-mass range $m_{\rm FP}/H_{0}\gtrsim 10^{32}$.
@@
The paper is organized as follows. In Section 2 we briefly review the bigravity framework and the dynamics of cosmological tensor modes in this theory. In Section 3, we study the dynamics of tensor modes in a de Sitter background. We obtain exact solutions for the mode functions of the mass eigenstates, for which we derive uniformly valid analytical approximations in terms of elementary functions. In Section 3.1 we identify different regimes of physical interest based on the propagation dynamics of the massive graviton. In each regime, we compute the mode functions in terms of redshift, the gravitational-wave luminosity distance, and study the propagation of wave packets. The coherence properties of gravitational radiation in the presence of a noisy environment, represented by classical incoherent matter sources, are studied in Section 4. We conclude in Section 5 by reviewing our results and discussing directions for future theoretical developments and possible applications of our results to tests of bimetric gravity using gravitational wave observations. The paper also includes two technical appendices: in Appendix A we detail the derivation of the asymptotic solution for the massive graviton mode, while in Appendix B we include details of the calculation of the correlation functions of the mass eigenstates presented in Section 4.
@@ 2 Review of the bigravity framework
where $R^{(g)}$ and $R^{(f)}$ are the Ricci scalars of the metrics $g_{ab}$ and $f_{ab}$, respectively. The constants $M_{g}$, $M_{f}$, and $m$ have physical dimensions of mass, whereas the constants $\beta_{n}$ are dimensionless. For later convenience, we define $\alpha\coloneqq M_{f}/M_{g}$. In particular, this ratio defines the GR-limit of the theory as $\alpha\to 0$. Finally, the $e_{n}(\mathbb{S})$ are the elementary symmetric polynomials of the matrix $\mathbb{S}$ [27, 28], with $\mathbb{S}$ defined in terms of the two metrics $g_{ab}$ and $f_{ab}$ as
@@
where $x^{i}$ are comoving coordinates, $\eta$ denotes conformal time, and $\gamma_{ij}$ is the three-dimensional Euclidean metric. The tensor perturbations $h_{ij}$ and $\tilde{h}_{ij}$ are transverse and traceless. As in GR, linear perturbations belonging to different irreducible representations of the rotation group do not couple to one another; hence, we will not consider scalar or vector perturbations. Moreover, we can decompose both tensor perturbations in a polarization basis ($h_{ij}=h_{+}e_{ij}^{+}+h_{\times}e_{ij}^{\times}$  and $\tilde{h}_{ij}=\tilde{h}_{+}e_{ij}^{+}+\tilde{h}_{\times}e_{ij}^{\times}$), and the bigravity field equations imply that only components with the same polarization are coupled. Henceforth, we will drop the polarization labels $+$ and $\times$. A given polarization component is then decomposed in Fourier modes, that is, $h=\frac{1}{(2\pi)^{3}}\int\mbox{d}\textbf{k}\,e^{i\textbf{k}\cdot\textbf{x}}h_{\textbf{k}}$  and $\tilde{h}=\frac{1}{(2\pi)^{3}}\int\mbox{d}\textbf{k}\,e^{i\textbf{k}\cdot\textbf{x}}\tilde{h}_{\textbf{k}}$. In momentum space, the linearized equations of motion for the Fourier components $h_{\textbf{k}}(\eta)$ and $\tilde{h}_{\textbf{k}}(\eta)$ consist of a coupled system of second-order differential equations for the two metric perturbations,
@@
where $\textbf{k}|$ k | is the wavenumber, a prime denotes derivative with respect to conformal time $\eta$, $y\coloneqq b/a$ is the ratio between the scale factors, and the coupling $\lambda(y)\coloneqq\beta_{1}y+\beta_{2}(1+c)y^{2}+\beta_{3}c\,y^{3}$ originates from the interaction terms in the bigravity action (2.1). Here, $\pi_{\textbf{k}}(\eta)$ is the Fourier transform of the anisotropic stress of matter. For a detailed derivation of the system (2.5), we refer the reader to Refs. [29, 30, 31]. Since matter fields only couple to $h_{\textbf{k}}$, this is the only tensor mode directly relevant for gravitational wave detection.
@@
$\tilde{\mu}_{\textbf{k}}^{\prime\prime}+\left(k^{2}-\frac{b^{\prime\prime}}{b}\right)\tilde{\mu}_{\textbf{k}}-\frac{c^{\prime}}{c}\left(\frac{\tilde{\mu}_{\textbf{k}}^{\prime}}{b}-\frac{b^{\prime}\tilde{\mu}_{\textbf{k}}}{b^{2}}\right)+\frac{m^{2}}{\alpha^{2}}ca^{2}y^{-2}\lambda(y)\left(\tilde{\mu}_{\textbf{k}}-y\,\mu_{\textbf{k}}\right)=0~{}.$ (2.6b)
@@ 3 Coupled tensor modes in a de Sitter background
with the mixing angle $\theta$ defined such that [33]
@@
In order to ensure the absence of ghost instabilities [34, 35], we impose the Higuchi bound $m_{\rm FP}^{2}\geq 2H^{2}$.
@@
where $J_{\xi}$ and $Y_{\xi}$ are the Bessel functions of the first and second kind, respectively. In these expressions, we introduced the shorthand notation $\xi\coloneqq\sqrt{9/4-(m_{\rm FP}/H)^{2}}$, and we made explicit the $k$-dependence of the integration constants. The overall $H^{-1}$ factor is just a convenient choice of normalization for the modes. We stress that the exact solutions derived above are fully general and valid for modes on all scales and for all values of the mass. However, for the sake of extracting concrete physical information on the behavior of the system, expression (3.7b) might be unwieldy. Thus, in the following, we analyze its asymptotics in a relevant regime of interest and provide an accurate approximation in terms of elementary functions. This considerably simplifies the analysis of deviations from GR and helps us better understand potential signatures of bimetric gravity in the propagation of tensor modes. The approximation derived below is also convenient in order to study the propagation of gravitational waves in real space, as analyzed in the following subsections.
@@
The above conditions on $\xi$ correspond to the mass range $13/4<m_{\rm FP}^{2}/H^{2}$, and we will work under this assumption in the remainder of the paper. We recall that the theoretical lower bound for the mass is smaller, as it is set by the Higuchi bound, $2\leq m_{\rm FP}^{2}/H^{2}$. However, the mass range not covered by the above approximation, $42\leq m_{\rm FP}^{2}/H^{2}\leq 13/$ 4, is a narrow interval and will be disregarded in the following.
@@ 3.1 Regimes of physical interest
This regime is defined by the condition $m_{\rm FP}/H\ll|\eta|k$ , which translates into $|\xi|\ll|\eta|k$. In this case, the approximate solution (3.8) obtained earlier for the $v$ mode can be further simplified to
@@
Hence, in terms of the redshift, the solutions for the tensor modes $h_{\textbf{k}}$ and $\tilde{h}_{\textbf{k}}$ in this small-mass regime are
@@
Next, we study the effect of mixing of mass eigenstates on the gravitational-wave luminosity distance. We compare the gravitational-wave luminosity distance in bigravity, associated with the tensor perturbation of the physical metric, $h_{\textbf{k}}(z)$, to the corresponding quantity in GR, $h^{\rm GR}_{\textbf{k}}(z)$. To make a meaningful comparison, both perturbations must be subject to the same initial conditions on the observable mode in the two theories. We impose $h_{\textbf{k}}(z_{0})=h^{\rm GR}_{\textbf{k}}(z_{0})=f_{\textbf{k}}$ at some redshift $z_{0}>0$ corresponding to signal emission, and also assume as initial condition that the perturbations of the second metric are initially unexcited $\tilde{h}_{\textbf{k}}(z_{0})=0$. For definiteness, we only consider positive frequency solutions, $h_{\textbf{k}}\sim e^{-ik\eta}$. Hence, using Eq. (3.14) and the above initial conditions, we obtain the following solutions in the two theories
@@
This result is valid in the regime where the massless and massive components of the signal cannot be temporally resolved at the detector (‘temporally unresolved’ regime). Eq. (3.17) is also consistent with the result obtained in Ref. [5] in the large $k$ limit using a different method (cf. Eq. (3.52) therein). The behavior of the luminosity distance in this regime is illustrated in Figure 1. In particular, we identify a mass threshold, corresponding to

### v2

@@ Abstract
We carry out a detailed analytical investigation of the propagation of gravitational waves in ghost-free bimetric gravity in a late-time de Sitter epoch. In this regime, the dynamical equations for the massless and massive graviton modes can be decoupled and solved exactly. We provide uniform approximations for the modes in terms of elementary functions, which are valid on all scales and for all viable mass windows. We identify different dynamical regimes for the system, depending on the propagation properties of the massive graviton, and whether the massless and massive components of the signal can be temporally resolved or not. In each regime, we compute the gravitational-wave luminosity distance as a function of redshift and study the propagation of wave packets. This allows for the derivation of a new observational bound for the ghost-free bimetric theory using the event GW170817. Further, by an explicit computation, we show that the massless and massive components of the signal retain their coherence also in the regime where they can be temporally resolved, even when couplings to incoherent matter degrees of freedom are included.
@@ 1 Introduction
A particularly relevant example is bimetric gravity [7, 8, 9], which extends GR by considering two independent, dynamical and nonlinearly interacting metric fields. The interactions between the two metrics are constructed in a specific way that ensures the absence of the Boulware–Deser ghost [8, 10]. To maintain ghost-freedom, matter fields are coupled to only one of the two metrics [11, 12]. The theory can also be interpreted as describing the nonlinear dynamics of two interacting spin-2 fields (gravitons), one massless and the other massive, carrying 7 dynamical degrees of freedom in total. The free parameters associated with the interaction terms in the bimetric action can be rearranged to a set of 5 physical parameters: the effective energy density of dark energy $\Omega_{\Lambda}$, the mixing angle $\theta$ between the massless and massive gravitons, the mass $m_{\rm FP}$ of the massive graviton, and two additional parameters that determine the screening mechanism on solar system scales [13, 14]. Bimetric gravity admits two important limiting cases: GR is recovered in the $\theta\to 0$ limit, whereas massive gravity is obtained in the limit $\theta\to\frac{\pi}{2}$. The theory is compatible with all observational tests performed so far [15, 16, 17, 18, 19, 20, 21, 14, 22, 23], and has the potential to alleviate cosmological tensions [24, 25, 26, 27].
@@
In bimetric gravity, tensor perturbations are produced in states that are linear superpositions of massless and massive gravitons. Therefore, as they propagate, they undergo a mixing phenomenon that displays some analogies with neutrino and neutral-kaon oscillations. This was first noticed in Ref. [28] for proportional backgrounds [^1] That is, backgrounds such that the two metrics are conformally related with constant conformal factor, $f_{ab}=C^{2}\,g_{ab}$. For a cosmological background, this is equivalent to having a constant ratio of the scale factors and same conformal times., and later revisited in Refs. [29, 30] for cosmological backgrounds. As discussed in Ref. [29], in the regime where the gravitational wave signal component carried by the massive graviton cannot be temporally resolved from the massless component, the presence of the massive graviton induces a modulation of the gravitational wave signal, referred to as ‘gravitational wave oscillations’. Moreover, for gravitational wave signals with limited extent (wave packets) and sufficiently large propagation distances, the difference in group velocities between the massless and massive modes leads to a separation of the corresponding wave packets [30]. This has two main effects: a constant suppression factor in the amplitude of the GR-like signal carried by the massless component, and the existence of a secondary signal, interpreted as a distorted ‘echo’, carried by the massive mode. Previous analyses have placed constraints on the theory using gravitational wave observations, excluding some regions of the $(\theta,m_{\mathrm{FP}})$-parameter space. In particular, for $m_{\mathrm{FP}}\gtrsim 10^{-23}\,\mathrm{eV}$, the mixing angle is restricted to $\theta\lesssim 10^{\circ}$. Specifically, the viable regions of parameter space, which are also compatible with all the other theoretical and observational constraints obtained so far [30, 13, 14, 22, 31], include (in units of the Hubble constant $H_{0}$): a low-mass range $1\lesssim m_{\rm FP}/H_{0}\lesssim 10^{4}$; an intermediate-mass range with $10^{6}\lesssim m_{\rm FP}/H_{0}\lesssim 10^{8}$ ; and a large-mass range $m_{\rm FP}/H_{0}\gtrsim 10^{32}$.
@@
The paper is organized as follows. In Section 2 we briefly review the bigravity framework and the dynamics of cosmological tensor modes in this theory. In Section 3, we study the dynamics of tensor modes in a de Sitter background. We obtain exact solutions for the mode functions of the mass eigenstates, for which we derive uniformly valid analytical approximations in terms of elementary functions. In Section 3.1 we identify different regimes of physical interest based on the propagation dynamics of the massive graviton. In each regime, we compute the mode functions in terms of redshift, the gravitational-wave luminosity distance, and study the propagation of wave packets. The coherence properties of gravitational radiation in the presence of a noisy environment, represented by classical incoherent matter sources, are studied in Section 4. We conclude in Section 5 by reviewing our results and discussing directions for future theoretical developments and possible applications of our results to tests of bimetric gravity using gravitational wave observations. The paper also includes two technical appendices: in Appendix A we detail the derivation of the asymptotic solution for the massive graviton mode, while in Appendix B we include details of the calculation of the correlation functions of the mass eigenstates presented in Section 4. We work with units such that the speed of light and the reduced Planck constant are both equal to one.
@@ 2 Review of the bigravity framework
where $R^{(g)}$ and $R^{(f)}$ are the Ricci scalars of the metrics $g_{ab}$ and $f_{ab}$, respectively. The positive constants $M_{g}$, $M_{f}$, and $m$ have physical dimensions of mass, whereas the constants $\beta_{n}$ are dimensionless. For later convenience, we define $\alpha\coloneqq M_{f}/M_{g}$. In particular, this ratio defines the GR-limit of the theory as $\alpha\to 0$. Finally, the $e_{n}(\mathbb{S})$ are the elementary symmetric polynomials of the matrix $\mathbb{S}$ [33, 34], with $\mathbb{S}$ defined in terms of the two metrics $g_{ab}$ and $f_{ab}$ as
@@
where $x^{i}$ are comoving coordinates, $\eta$ denotes conformal time, and $\gamma_{ij}$ is the three-dimensional Euclidean metric. The tensor perturbations $h_{ij}$ and $\tilde{h}_{ij}$ are transverse and traceless. As in GR, linear perturbations belonging to different irreducible representations of the rotation group do not couple to one another; hence, we will not consider scalar or vector perturbations. [^2] We note, however, that on cosmological backgrounds the scalar sector of bimetric gravity is typically plagued by a gradient instability at early times. The existence of this instability suggests a possible breakdown of linear perturbation theory rather than a physical instability, and developing a consistent framework to describe the evolution of small perturbations on cosmological backgrounds remains an open problem. For works on the instability of the scalar sector, see Refs. [35, 36, 37, 38, 39, 40, 41, 19]. In this work, we focus on linear tensor perturbations, which are unaffected by this instability. We can then decompose both tensor perturbations in a polarization basis ($h_{ij}=h_{+}e_{ij}^{+}+h_{\times}e_{ij}^{\times}$  and $\tilde{h}_{ij}=\tilde{h}_{+}e_{ij}^{+}+\tilde{h}_{\times}e_{ij}^{\times}$), and the bigravity field equations imply that only components with the same polarization are coupled. Henceforth, we will drop the polarization labels ++ and $\times$. A given polarization component is then decomposed in Fourier modes, that is, $h=\frac{1}{(2\pi)^{3}}\int\mbox{d}\textbf{k}\,e^{i\textbf{k}\cdot\textbf{x}}h_{\textbf{k}}$  and $\tilde{h}=\frac{1}{(2\pi)^{3}}\int\mbox{d}\textbf{k}\,e^{i\textbf{k}\cdot\textbf{x}}\tilde{h}_{\textbf{k}}$. In momentum space, the linearized equations of motion for the Fourier components $h_{\textbf{k}}(\eta)$ and $\tilde{h}_{\textbf{k}}(\eta)$ consist of a coupled system of second-order differential equations for the two metric perturbations,
@@
where $k=|\textbf{k}|$ is the wavenumber, a prime denotes derivative with respect to conformal time $\eta$, $y\coloneqq b/a$ is the ratio between the scale factors, and the coupling $\lambda(y)\coloneqq\beta_{1}y+\beta_{2}(1+c)y^{2}+\beta_{3}c\,y^{3}$ originates from the interaction terms in the bigravity action (2.1). Here, $\pi_{\textbf{k}}(\eta)$ is the Fourier transform of the anisotropic stress of matter. We note that these equations hold for any general perturbed cosmological background of the form (2.4) (that is, regardless of the specific functional form of $a(\eta)$, $b(\eta)$, $c(\eta)$). For a detailed derivation of the system (2.5), we refer the reader to Refs. [42, 43, 44]. Since matter fields only couple to $h_{\textbf{k}}$, this is the only tensor mode directly relevant for gravitational wave detection.
@@
$\tilde{\mu}_{\textbf{k}}^{\prime\prime}+\left(c^{2}k^{2}-\frac{b^{\prime\prime}}{b}\right)\tilde{\mu}_{\textbf{k}}-\frac{c^{\prime}}{c}\left(\tilde{\mu}_{\textbf{k}}^{\prime}-\frac{b^{\prime}\tilde{\mu}_{\textbf{k}}}{b}\right)+\frac{m^{2}}{\alpha^{2}}ca^{2}y^{-2}\lambda(y)\left(\tilde{\mu}_{\textbf{k}}-y\,\mu_{\textbf{k}}\right)=0~.$ (2.6b)
@@ 3 Coupled tensor modes in a de Sitter background
with the mixing angle $\theta\in(0,\pi/2)$ defined such that [^3] Note that this is slightly different from the definition adopted in Ref. [20], where the mixing angle is defined as $\alpha y_{*}$.
@@
In order to ensure the absence of ghost instabilities [46, 47], we impose the Higuchi bound $m_{\rm FP}^{2}>2H^{2}$.
@@
where $A_{\textbf{k}},\,\varphi_{\textbf{k}},\,C_{\textbf{k}}$ and $D_{\textbf{k}}$ are integration constants and $J_{\xi}$ and $Y_{\xi}$ are the Bessel functions of the first and second kind, respectively. In these expressions, we introduced the shorthand notation $\xi\coloneqq\sqrt{9/4-(m_{\rm FP}/H)^{2}}$, and we made explicit the $k$-dependence of the integration constants. The overall $H^{-1}$ factor is just a convenient choice of normalization for the modes. We stress that the exact solutions derived above are fully general and valid for modes on all scales and for all values of the mass. This result goes beyond previous studies, which are restricted to sub-horizon scales [29, 30, 32]. However, for the sake of extracting concrete physical information on the behavior of the system, expression (3.7b) might be unwieldy. Thus, in the following, we analyze its asymptotics in a relevant regime of interest and provide an accurate approximation in terms of elementary functions. This considerably simplifies the analysis of deviations from GR and helps us better understand potential signatures of bimetric gravity in the propagation of tensor modes. The approximation derived below is also convenient in order to study the propagation of gravitational waves in real space, as analyzed in the following subsections.
@@
The above conditions on $\xi$ correspond to the mass range $13/4<m_{\rm FP}^{2}/H^{2}$, and we will work under this assumption in the remainder of the paper. We recall that the theoretical lower bound for the mass is smaller, as it is set by the Higuchi bound, $2<m_{\rm FP}^{2}/H^{2}$. However, the mass range not covered by the above approximation, $2<m_{\rm FP}^{2}/H^{2}\leq 13/4$, is a narrow interval and will be disregarded in the following.
@@ 3.1 Regimes of physical interest
This regime is defined by the condition $m_{\rm FP}/H\ll|\eta|k$ , which translates into $|\xi|\ll|\eta|k$. Physically, it corresponds to the sub-horizon limit (i.e., $1\ll|\eta|k$), due to the Higuchi bound on the mass, $2<m_{\rm FP}^{2}/H^{2}$. Correspondingly, the physical wavelength is much shorter than the length scale $1/m_{\rm FP}$ set by the graviton mass.
In this case, the approximate solution (3.8) obtained earlier for the $v$ mode can be further simplified to
@@
Hence, in terms of the redshift, the solutions for the tensor modes $h_{\textbf{k}}$ and $\tilde{h}_{\textbf{k}}$ in this regime are
@@
Next, we study the effect of mixing of mass eigenstates on the gravitational-wave luminosity distance. We compare the gravitational-wave luminosity distance in bigravity, associated with the tensor perturbation of the physical metric, $h_{\textbf{k}}(z)$, to the corresponding quantity in GR, $h^{\rm GR}_{\textbf{k}}(z)$. To make a meaningful comparison, in the two theories perturbations must be subject to the same initial conditions on the observable mode. We impose $h_{\textbf{k}}(z_{0})=h^{\rm GR}_{\textbf{k}}(z_{0})=f_{\textbf{k}}$ at some redshift $z_{0}>0$ corresponding to signal emission, and also assume as initial condition that the perturbations of the second metric are initially unexcited $\tilde{h}_{\textbf{k}}(z_{0})=0$. This is expected due to the screening mechanism and the fact that only one metric couples to matter [30]. For definiteness, we only consider positive frequency solutions, $h_{\textbf{k}}\sim e^{-ik\eta}$. Hence, using Eq. (3.14) and the above initial conditions, we obtain the following solutions in the two theories
@@
This expression is consistent with the result obtained in Ref. [5] in the large $k$ limit using a different method (cf. Eq. (3.52) therein). The behavior of the luminosity distance in this regime is illustrated in Figure 1. In particular, we identify a mass threshold, corresponding to
@@
For sufficiently large values of the mass, such that we can approximate $|\xi|^{2}\approx m_{\rm FP}^{2}/H^{2}$, the threshold (3.18) corresponds to
$\left(\frac{m_{\rm FP}}{10^{-22}\,{\rm eV}}\right)^{2}\approx 8.8\times 10^{-2}\,h\left(\frac{k}{100\,{\rm Hz}}\right)~,$ (3.19)
where $h$ is the dimensionless Hubble rate, defined according to $H=100\,h\,{\rm km}\,{\rm s}^{-1}{\rm Mpc}^{-1}$. The behavior of the luminosity distance varies significantly depending on whether the mass lies above or below this threshold. Specifically, on the one hand, for values of the mass such that $|\xi|^{2}\leq 2\pi k/H$, the ratio of luminosity distances (3.17) is a monotonically increasing function of the redshift $z_{0}$ and, for large redshift values, it approaches the limit
$\lim_{z_{0}\to+\infty}\frac{d_{L}^{\text{gw}}(z_{0})}{d_{L}^{\text{em}}(z_{0})}=\frac{1}{\cos^{2}\theta}\left[1+\tan^{4}\theta+2\tan^{2}\theta\cos\left(\frac{|\xi|^{2}H}{2k}\right)\right]^{-1/2}~.$ (3.20)
On the other hand, for $|\xi|^{2}>2\pi k/H$, (3.17) transitions to an oscillatory behavior, subject to the bounds
$1\leq d_{L}^{\text{gw}}(z_{0})/d_{L}^{\text{em}}(z_{0})\leq|\cos(2\theta)|^{-1}~.$ (3.21)
For $\theta\neq\frac{\pi}{4}$, the total number of peaks and troughs is $\left\lceil\frac{H|\xi|^{2}}{2\pi k}-1\right\rceil$. Interestingly, the upper bound is divergent when the mixing angle is $\theta=\frac{\pi}{4}$ (maximal mixing). Physically, such zeroes stem from the fact that, in the regime at hand and for $\theta=\pi/4$ , Eq. (3.15b) represents a superposition of two equal-amplitude waves with a small $z_{0}$-dependent beat frequency. While the GR amplitude $|h^{\rm GR}_{\textbf{k}}(z=0)|$ is always nonzero, the amplitude of the bigravity signal $|h_{\textbf{k}}(z=0)|$ vanishes at some $z_{0}>0$ where the massive and massless signal components interfere destructively. Hence, this divergence simply reflects complete destructive interference in the maximally mixed case.
We remark that for sub-horizon modes, $|\xi|\ll|\eta|k$, the threshold (3.18) lies well within the accessible range of $|\xi|$ and is therefore physically relevant. The existence of two different regimes can be understood intuitively as follows. Below the threshold (3.18), the relative phase between the massless and massive components is a monotonically increasing function of redshift and strictly less than $\pi$, even for an infinitely distant source (in which case it tends asymptotically to a constant). Hence, even though destructive interference between the two components implies an increase of $d_{L}^{\text{gw}}(z_{0})/d_{L}^{\text{em}}(z_{0})$ with redshift, the behavior is monotonic and bounded from above. On the other hand, above the threshold (3.18), the relative phase exceeds $\pi$, leading to a succession of phases where $d_{L}^{\text{gw}}(z_{0})/d_{L}^{\text{em}}(z_{0})$ is monotonically increasing or decreasing. We note that the existence of the threshold (3.18) itself is a peculiarity of the nonlinear dependence of the massive mode on the redshift.
Finally, we observe that the luminosity distance (3.17) cannot be modelled using the phenomenological parametrization proposed in Ref. [4], which assumes
$\left(\frac{d_{L}^{\text{gw}}(z_{0})}{d_{L}^{\text{em}}(z_{0})}\right)^{\rm pheno}=\Xi_{0}+\frac{1-\Xi_{0}}{(1+z_{0})^{n}}~,$ (3.22)
with free parameters $\Xi_{0}$, $n$. The two formulas (3.17) and (3.22) are clearly incompatible in the above-threshold regime, due to the oscillatory nature of the bigravity luminosity distance. They are also incompatible in the below-threshold regime: even though at high redshift both expressions asymptote to a constant value, and both are monotonically increasing for $n>0$ and $\Xi_{0}>1$, their functional forms are clearly different. Moreover, they predict a different behavior of the luminosity distance at low redshift, where we have the following asymptotics
$\frac{d_{L}^{\text{gw}}(z_{0})}{d_{L}^{\text{em}}(z_{0})}\approx 1+\frac{H^{2}|\xi|^{4}}{32k^{2}}\sin^{2}(2\theta)z_{0}^{2}+{\cal O}(z_{0})^{3}~,$~ (3.23a)
$\left(\frac{d_{L}^{\text{gw}}(z_{0})}{d_{L}^{\text{em}}(z_{0})}\right)^{\rm pheno}\approx 1+n(\Xi_{0}-1)z_{0}+{\cal O}(z_{0})^{2}~,$ (3.23b)
which are incompatible to leading order in $z_{0}$. Moreover, we note that $d_{L}^{\text{gw}}$ in bigravity depends on the wave-number $k$, whereas the phenomenological model (3.22) assumes that $\Xi_{0}$, $n$ do not dependent on $k$. We conclude that observational constraints based on the phenomenological parametrization (3.22) cannot be translated into constraints on the parameter space of bigravity.
The result (3.17) is valid in the regime where the massless and massive components of the signal cannot be temporally resolved at the detector (‘temporally unresolved’ regime), since the observed signal is described by the full expression in Eq. (3.15b). In turn, in the regime where the massless and massive components can be resolved at the detector (‘temporally resolved’ regime), the luminosity distance is determined only by the massless component, represented by the first term in the solution (3.15b). (As explained later in this section, the massive component represents a delayed ‘echo’ of the original signal.) The amplitude of the massless component is suppressed by a factor $\cos^{2}\theta$ compared to the corresponding solution in GR. Hence, the luminosity distance in the ‘temporally resolved’ regime reads
Multimessenger events, based on the combined observation of gravitational-wave and electromagnetic signals emitted from the same astrophysical source, can be used to constrain the bigravity parameter space $(\theta,m_{\rm FP})$. For illustrative purposes, we focus on the low-redshift event GW170817 [50] observed by LIGO-Virgo detectors and its electromagnetic counterpart. We start by observing that, for sufficiently large values of $m_{\rm FP}$, we can approximate $|\xi|\approx m_{\rm FP}/H$ in Eq. (3.23a), which can then be recast as
$m^{2}_{\rm FP}\sin(2\theta)=\frac{4\sqrt{2}\,Hk}{z_{0}}\sqrt{\frac{d_{L}^{\text{gw}}}{d_{L}^{\text{em}}}-1}~.$ (3.25)
For GW170817, we have $d_{L}^{\text{gw}}=43.8^{+2.9}_{-6.9}\,{\rm Mpc}$, $d_{L}^{\text{em}}=40.7\pm 1.4\,{\rm(stat)}\pm 1.9\,{\rm(sys)}\,{\rm Mpc}$, $z_{0}\simeq 0.0098$ [4, 51, 52, 53]. We consider the minimal frequency of the signal, which corresponds to $k=2\pi\times 24\,{\rm Hz}$ [50]. Using Eq. (3.25) with these experimental data, and combining the experimental errors in quadrature, [^4] A similar approach was followed in Refs. [4, 51] to derive constraints on the phenomenological model (3.22). we obtain the constraint $(m_{\rm FP}^{2}/H)\sin(2\theta)<2.6\times 10^{-11}\,{\rm eV}$. This can be re-expressed in terms of the reduced Hubble rate $h$ as
$\left(\frac{m_{\rm FP}}{10^{-22}\,{\rm eV}}\right)^{2}\sin(2\theta)<5.5\,h~.$ (3.26)
The constraint (3.26) still allows for significant deviations of $d_{L}^{\text{gw}}/d_{L}^{\text{em}}$ from GR at large redshift. Importantly, it is also much milder compared to bounds on $m_{\rm FP}$ in massive gravity, due to the extra freedom represented by the mixing angle. An exclusion plot illustrating the constraint (3.26) in the $(\theta,m_{\mathrm{FP}})$ plane is shown in Figure 2. We recall that this bound has been derived under the assumption that the source distance is such that the massless and massive signal components cannot be temporally resolved at the detector. Further constraints based on the non-observation of waveform distortions in the ‘temporally unresolved’ regime, as well as amplitude suppression and echoes in the ‘temporally resolved’ regime, have been analyzed in Refs. [29, 30, 31] (where such regimes are referred to as ‘coherent’ and ‘decoherent’, respectively). These are complementary to the bound (3.26).